An oil well cost $2,030,000 and is calculated to hold 350,000 barrels of oil. There is no residual value. Which journal entryFleetwood Company signed a one-year note payable for $58,000 at 8% annual interest. What is the interest expense for 2019 ifSales revenue for a sporting goods store amounted to $535,000 for the current period. All sales are on account and are subjecCarpenters Company, a manufacturing company, acquired equipment on January 1, 2017 for $570,000. Estimated useful life of theAn asset was purchased for $37,000 on January 1, 2019. The assets estimated useful life was five years, and its residual val
Answer #1

1) Journal entry

Date account and explanation Debit Credit
Depletion expense (2030000/350000*47000) 272600
Accumulated depletion-Oil reserve 272600

So answer is c)

2) Interest expense = 58000*8%*5/12 = 1933

So answer is b) $1933

3) Journal entry

Date account and explanation Debit Credit
Account receivable (535000*1.1) 588500
Sales revenue 535000
Sales tax payable 53500

So answer is b) a debit to account receivable $588500

4) Accumulated depreciation = 570000-10000/7*3 = 240000

2020 Depreciation expense = (570000-240000-10000)/6 = 53333

So answer is d) $53333

5) Depreciation expense = (37000-8000/5) = 5800

Book value = 37000-5800 = 31200

Gain (loss) = 23000-31200 = -8200

So answer is b) $8200 loss